CITATIONS

James T. Bonnen, 1968. “The Distribution of Benefits from Selected U.S. Farm Programs,” in the President’s National Advisory Committee on Rural Poverty, Rural Poverty in the United States, Washington, D.C., May.

Phelim Boyle, 1989. “The Quality Option and Timing Option in Futures

Contracts,” Journal of Finance, Vol. 44, pp. 101–13. Chicago Board of Trade, 1989. 1988 Wheat Study; 1988 Corn Study; 1988

Soybean Study, Vols. I and II, Economic Analysis and Planning Department, Chicago.

General Accounting Office (GAO), 1991. Chicago Futures Markets: Selecting Agricultural Futures Delivery Points Involves Tradeoffs, Washington, D.C., June.

Roger W. Gray and Anne E. Peck, 1981. “The Chicago Wheat Futures Market: Recent Problems in Historical Perspective,” Food Research Institute Studies, Vol. 18, pp. 89–115.

Thomas A. Hieronymus, 1977. The Economics of Futures Trading for Commercial and Personal Profit, 2nd ed., Commodity Research Bureau, New Jersey.

George W. Hoffman, 1932. Future Trading Upon Organized Commodity

Markets in the United States, University of Pennsylvania, Philadelphia. Harold S. Irwin, 1954. Evolution of Futures Markets, Mimir Publishers,

Madison. Kandice Kahl, 1983. “Determination of the Recommended Hedging Ratio,”

American Journal of Agricultural Economics, Vol. 65, pp. 603–05. Jan Kmenta, 1986. Elements of Econometrics, 2nd ed., Macmillan, New

York. Mid-America Institute for Public Policy, 1991. Grain Futures Contracts for

the 1990s: An Economic Appraisal, Chicago Board of Trade, Chicago, July.

Allen Paul, 1970. “Pricing of Binspace: A Contribution to the Theory of Storage,” American Journal of Agricultural Economics, Vol. 52, pp. 1–12.

Allen Paul, Kandice Kahl, and William Tomek, 1981. Performance of Futures Markets: The Case of Potatoes, U.S. Department of Agriculture,

Food Research Institute Studies, Vol. XXII, No. 2, 1991.

Technical Bulletin 1636, Washington, D.C. Anne E. Peck and Jeffrey C. Williams, 1991. An Evaluation of the Perfor-

mance of the Chicago Board of Trade Wheat, Corn, and Soybean Futures Contracts During Delivery Periods from 1964–65 Through 1988–89, National Grain and Feed Association, Washington, D.C., April.

Richard Sandor, 1973. “Innovation by an Exchange: A Case Study of the Development of the Plywood Futures Contract,” Journal of Law and Economics, Vol. 16, pp. 119–36, reprinted in Anne E. Peck, ed., Views from the Trade, Chicago Board of Trade, Chicago, 1978.

Roger Silk, 1988. “Implicit Delivery Options in Futures Contracts and Optimal Exercise Strategy: Theory and Evidence,” Ph.D. dissertation, Stanford University, Stanford.

Sarahelen Thompson, Robert J. Hauser, and James S. Eales, 1990. “An Empirical Analysis of Cash and Futures Grain Price Relationships in the North Central Region,” North Central Journal of Agricultural Economics, Vol. 12, pp. 241–54.

Jeffrey C. Williams and Anne E. Peck, 1991. “The Timing Option in Com-

modity Futures Contracts,” Stanford University, February. Holbrook Working, 1948. “The Theory of the Inverse Carrying Charge,”

Journal of Farm Economics, Vol. 30, pp. 1–28, reprinted in Anne E. Peck, ed., Selected Writings of Holbrook Working, Chicago Board of Trade, Chicago, 1977.

, 1949. “The Theory of Price of Storage,” American Economic Review, Vol. 39, pp. 1254–63, reprinted in Anne E. Peck, Selected Writings of Holbrook Working, Chicago Board of Trade, Chicago, 1977.

, 1953. “Hedging Reconsidered,” Journal of Farm Economics, Vol. 35, pp. 544-61, reprinted in Anne E. Peck, ed., Selected Writings of Holbrook Working, Chicago Board of Trade, Chicago, 1977.